# Communication Constellation Design of Minimum Number of Satellites with Continuous Coverage and Inter-Satellite Link

### v1

@@ 3.1 Geometry of Earth Coverage
Figure 3a describes the geometric relationships between the satellite, coverage edge, and center of the Earth are advantageous for coverage analysis. The true horizon is tangent from the satellite to the Earth’s surface. The angle between the true horizon and the subsatellite point is called the maximum Earth central angle (ECA, $\lambda_{0}$) or the angular radius of the Earth ($\rho$) when measured from the center of the Earth and the satellite, respectively. When the satellite is located at an altitude of h km, the angular radius of the Earth is determined by Eq. (32), where $R_{E}$ is the Earth’s radius.
@@
Usually, the payload’s coverage determines the coverage performance of a single satellite that constitutes the constellation. The Earth central angle ($\lambda$) measures the size of the payload’s coverage ($\eta$) on the surface of the Earth and is defined as the angular distance between the subsatellite point (SSP) and the coverage edge. The coverage edge is the rim of the satellite’s coverage on Earth. The elevation angle ($\varepsilon$) is measured at the coverage edge from the local horizontal to the satellite.
The coverage can be approached from two perspectives: the satellite and the target area (Figure 3b). When considering the satellite perspective, it’s important to evaluate the payload’s specification. According to the definition of the coverage, the nadir angle ($\eta=\eta(t)$) should be smaller than the payload’s beam coverage ($\eta_{Max}$). From the target point perspective, it is considered as covered when the elevation angle of the satellite ($\varepsilon=\varepsilon(t)$) is greater than the target point’s minimum elevation angle ($\varepsilon_{min}$).
The trigonometry of the blue shaded triangle in Figure 3a derives the formulae between $\eta$, $\varepsilon$, and $\lambda$ as Eqs. (33) and (34).
@@
For the communication constellations, the constraints are imposed on both the payload specification and the elevation angle of the target point. Therefore, the trigonometry in Eqs. (33) and (34) is crucial as it prevents redundant computational complexity. For example, suppose that the beam coverage of the spacecraft ($\eta_{Max}$) is $45$ deg and the ground station has a minimum elevation angle ($\varepsilon_{min}$) of $30$ deg. If the satellite’s altitude is $1,200$ km, the angular radius of the Earth ($\rho$) is $57.31$ deg. The Eq. (33) immediately converts the payload’s coverage to the elevation angle as
@@
where $\overline{\varepsilon}$ is the elevation angle that corresponds to $\eta_{Max}$ and does not have a physical meaning. The ECA ($\overline{\lambda}$) is calculated from Eq. (34) as
@@
In the same manner, the minimum elevation angle ($\varepsilon_{min}$) derives its corresponding parameters as $\tilde{\eta}=46.79$ deg and $\tilde{\lambda}=13.21$ deg. The ECA is the visualized size of the coverage on the Earth’s surface and the smaller the ECA is, the more degraded the coverage performance is. Therefore, the simulation only with $\overline{\lambda}$, $\eta_{Max}$, or $\overline{\varepsilon}$ is enough to analyze if the constellation satisfies the coverage requirement. Since the coverage analyses with $\overline{\lambda}$, $\eta_{Max}$, or $\overline{\varepsilon}$ show the same results but only conducted at different perspectives, the simulation with only one of three constraints reduces the computational cost.
@@ 3.2 Voronoi Tessellation
The problem of continuous coverage constellation can be reduced to obtaining the circumradius of the adjacent three points. For a set of discretized points, the circumradius of the adjacent three points can be defined. When the circumradius is less than a specified value, the distance from any point within the region is shorter than the specified value. References 2 and 3 introduced the satellite triad method as the coverage analysis technique for the Walker constellation. The research subject of the references was the constellation of less than 20 satellites. This research utilizes the Delaunay triangulation method that is first suggested in the reference 4 to generalize the number of satellites of the constellation.
The Delaunay triangulation is a computational geometry subdividing the discretized points into triangles[9]. This algorithm defines the Delaunay criterion to construct the Delaunay conformant triangles that do not contain other points inside the circumcircles. The Voronoi diagram (VD) is the dual graph of the Delaunay triangle (DT) and is drawn by connecting the circumcenters of the Delaunay triangles. The Voronoi tessellation is the tiling of a plane or a sphere with Voronoi diagrams. If the tiled region is a restricted closed area on the sphere, the Voronoi diagrams have boundaries cut by the boundary and are called bounded Voronoi diagram (BVD) [10, 4]. The constellation coverage problem can be formulated as the Voronoi tessellation problem. The subsatellite points on the Earth’s surface are the discretized points on the three-dimensional spherical surface. The area of interest is not the whole globe, the Voronoi diagram is bounded within the target region. The spherical Delaunay triangles and spherical bounded Voronoi diagram will derive the solution, but the word ’spherical’ will be omitted for brevity from here on.
@@
Figure 4 depicts an example of the Delaunay triangle, Voronoi diagram, and the bounded Voronoi diagram. In Figure 4(a), there are six DTs that contains $p_{k}$ as their vertices for an arbitrary subsatellite point $p_{k}$, where $k=1,2,...,T$. The ${DT}_{k,l}$ are depicted as red triangles, and the circumcenters $C_{k,l}$ are blue dots, where $l=1,2,...,N_{k}$. The $N_{k}$ is the number of triangles that have the vertice $p_{k}$, and can be different for each $p_{k}$. Any subsatellite point posseses only one VD, and the $VD_{k}$ is drawn by connecting the blue dots, the circumcenters $\textbf{C}_{k}=\left\{C_{k,l}\mid l=1,2,...,N_{k}\right\}$, as in Figure 4(b).
The Voronoi diagram in Figure 4(b) can be utilized for the global coverage problem. However, for the regional coverage problem, the Voronoi diagram should be bounded within the area of interest. Especially, for the regional continuous coverage problem, the Voronoi diagram bounded within the latitude range of AoI as a circular band can efficiently analyze the continuity of the constellation[4]. Suppose that the northern area of interest, then the bounded Voronoi diagram appears Figure 4(c). Then, $p_{k}$ has ${BVD}_{k}$ with different vertices $\overline{\textbf{C}}_{k}=\left\{\overline{C}_{k,\overline{l}}\mid\overline{l}=1,2,...,\overline{N}_{k}\right\}$.
Let us define the angular distance $\theta_{k,\overline{l}}$ as the angular distance between the subsatellite point $p_{k}$ and the vertices $\overline{C}_{k,\overline{l}}$. Then, the maximum distance of the ${BVD}_{k}$ of the $k$th satellite is expressed as
@@
As a result, the maximum angular distance of the whole constellation is obtained as
@@
Assuming the homogeneous constellation, the coverage performances of all the satellites are the same and can be expressed as $\lambda^{*}$. Then, the continuous coverage problem statement is defined as
@@
For the global coverage problem, the maximum angular distance should be defined from $\theta_{k,l}$ and $C_{k,l}$, and the rest procedures are the same.
@@ 3.3 APC Decomposition
The reference 7 developed the APC decomposition from the circular convolution phenomenon between a seed satellite’s access profile, a constellation pattern vector, and a coverage timeline. The access profile between the $k$th satellite and the $j$th target point ($\textbf{{v}}_{k,j}$) defines its elements as
@@
where $\tau$ is the discretized time variable, and $\mathcal{J}$ is the set of target points.
The coverage timeline of the constellation $\textbf{{b}}_{j}$ is derived as the summation of all the access profiles:
@@
The circular convolution phenomenon expresses the coverage timeline $\textbf{{b}}_{j}$ with respect to the seed-satellite access profile $\textbf{{v}}_{0,j}$ and the coverage pattern vector x as
@@
This circular convolution operation can be described in a linearized form as
@@
In summary, the coverage timeline of the whole constellation is obtained from the circular convolution of the seed-satellite access profile and the constellation pattern vector. This concept of APC decomposition reduces the optimal CGT constellation design problem to the constellation pattern vector optimization problem.
@@ 4 Relative Motion in Adjacent Orbital Planes
Since the Walker-Delta constellation satellites are designed to have the same altitude and inclination, the relative motion between ${SAT}_{m+1,n}$ and ${SAT}_{m,n}$ can be described analytically [11]. The minimum and maximum relative angular distances ($\theta_{min}$ and $\theta_{max}$) are formulated as
@@
The spherical triangle in Figure 5 derives the geometric relationship between the relative phase $\phi_{R}$ and angular distances $\phi_{m}$ and $\phi_{m+1}$ as
@@
where the differential angular distance $\phi_{m+1}-\phi_{m}$ is the relative argument of latitude $\Delta u$ in the Walker-Delta constellation. The formula to calculate $\phi_{m+1}$ is as follows:
@@
As a result, Eq. (51) provides the argument of $\phi_{R}$ as
@@
The inter-satellite link (ISL) constraints the range of relative motion so that the signal is not interfered within the link margin. Therefore, the minimum and maximum relative distances in Eqs. (46) and (47) contained within the specified range could guarantee smooth ISL communication.
The relative motion in the adjacent orbital plane described in Eqs. (48), (50), and (51) explains the relative motion between ${SAT}_{m,n}$ and ${SAT}_{m+1,n}$ where $m=1,...,P-1$ and $n=1,...,S$. When $m=P$, the relative motion between ${SAT}_{P,n}$ and ${SAT}_{1,n+F}$ is proven the same with the one between ${SAT}_{m,n}$ and ${SAT}_{m+1,n}$ by Eqs. (1), (2), (48), (50) and (51). As a result, if Eqs. (46) and (47) satisfy the constraint on the ISL link, then all the ISL links are connected without any isolated link.
@@ 5.1 Continuous Coverage Analysis
The constellation is designed with the three constellation design methods introduced in the previous sections: Walker-Delta constellation, Quasi-symmetric CGT constellation, and BILP CGT constellation. The seed satellite for the three constellations is designed to have the repetition period $\nu=14/1$. The inclination is 42 degrees which is 3–5 degrees higher than the area of interest[12, 13]. The minimum elevation angle $\varepsilon_{min}$ is 15 degrees and the target point is located at Seoul. The Walker-Delta constellation is analyzed assuming two-body motion and the target area is the circle of radius 100 km around Seoul. The CGT constellations assume $J_{2}$ perturbation and a single target point. The sampling time or time step $t_{step}$ is 1 and 300 seconds for Walker-Delta and CGT constellations respectively, and the simulation time horizon is 1 day for both.
@@ 5.1.1 Walker-Delta Constellation
Figure 6 depicts the bounded Voronoi diagram simulation result of Walker-Delta constellation. The global search from the less number of satellites obtains that 40 satellites at the inclination 42 $\deg$ is the minimum number of satellites to achieve the continuous coverage by Walker-Delta constellation. The empty circles are the $T/P/F$ parameters that have duplicate positions in Eq. (14) and are preclude from coverage analysis. The empty diamond markers mean that the parameters could not achieve continuous coverage. The continuous coverage solution is $i$: $T/P(S)/F$ = $42$: $40/40(1)/30$ and denoted as the black diamond. The ECA of this solution ($\lambda^{*}$) is $15.86\deg$.
5.1.2 Common Ground-track Constellations
@@
The CGT constellation’s pattern reveals its characteristics in the $\left(\Omega,M\right)$ space such as the period ratio and symmetricity (Figure 7). The gradient of admissible set is $-N_{P}/N_{D}$ and equal to $-\nu$ by Eqs. (25) and (15). The constellation pattern vectors are laid on some points along the admissible set. The quasi-symmetric set configures symmetrically in Figure 7 as $\textbf{{x}}_{qs}$ on the second panel of Figure 8a is equally spaced. Because the $N_{qs}$ is 32 and $L$ is 288, $L/N_{qs}$ is divided as the integer 9 and the spacing is perfectly symmetric. On the other hand, the BILP constellation pattern vector is irregularly spaced in Figures 7 and 8b. However, the BILP constellation achieves less number of satellites $N_{bilp}$ as 31 while both constellations exhibit the single-fold coverage.
@@ 5.2.1 Walker-Delta Constellation
The minimum and maximum relative distances are the line distance calculated from the angular distances in Eqs. (46) and (47) (Figure 9). The upper and lower bound of bars reveal the relative distance ranges in the adjacent orbital planes. The color of the bars distinguish the number of planes and the x axis is $F$ numbers. Since this graph shows the relative motion at a glance, it is a useful tool for ISL connectivity analysis. The continuous coverage solution $i$: $T/P(S)/F$ = $42$: $40/40(1)/30$ has the relative motion range from $9559.77km$ to $9589.64km$.

### v2

@@ 3.1 Geometry of Earth Coverage
Figure 3a shows the geometric relationships between the satellite, coverage edge, and center of the Earth and is advantageous for coverage analysis. The true horizon is tangential from the satellite to the Earth’s surface. The angle between the true horizon and the subsatellite point (SSP) is called the maximum Earth central angle (ECA, $\lambda_{0}$) or the angular radius of the Earth ($\rho$) when measured from the center of the Earth and the satellite, respectively. When the satellite is located at an altitude of h km, the angular radius of the Earth is determined using Eq. (32):
@@
where $R_{E}$ is the Earth’s radius.
Usually, the payload’s coverage determines the coverage performance of a single satellite that constitutes the constellation. The Earth central angle ($\lambda$) measures the size of payload coverage ($\eta$) on the Earth’s surface and is defined as the angular distance between the SSP and the coverage edge. The coverage edge is the rim of the satellite coverage on the Earth. The elevation angle ($\varepsilon$) is measured at the coverage edge from the local horizontal to the satellite.
The coverage can be approached from two perspectives: the satellite and the target area (Figure 3b). When considering the satellite perspective, it is important to evaluate payload’s specifications. According to the definition of the coverage, the nadir angle ($\eta=\eta(t)$) should be smaller than the payload beam coverage ($\eta_{Max}$). From the target point perspective, it is considered to be covered when the elevation angle of the satellite ($\varepsilon=\varepsilon(t)$) is greater than the target point’s minimum elevation angle ($\varepsilon_{min}$).
The trigonometry of the blue shaded triangle in Figure 3a yields the formulae for $\eta$, $\varepsilon$, and $\lambda$ as Eqs. (33) and (34).
@@
For communication constellations, constraints are imposed on both the payload specification and the elevation angle of the target point. Therefore, the trigonometry of Eqs. (33) and (34) is crucial, as it prevents redundant computational complexity. For example, suppose that the beam coverage of the spacecraft ($\eta_{Max}$) is $45\deg$, and the ground station has a minimum elevation angle ($\varepsilon_{min}$) of $30\deg$. If the satellite’s altitude is $1,200$ km, the angular radius of Earth ($\rho$) is $57.31\deg$. The Eq. (33) immediately converts the payload’s coverage to the elevation angle as
@@
where $\overline{\varepsilon}$ is the elevation angle corresponding to $\eta_{Max}$ and does not have a physical meaning. The ECA ($\overline{\lambda}$) is calculated using Eq. (34) as
@@
In the same manner, the minimum elevation angle ($\varepsilon_{min}$) yields its corresponding parameters as $\tilde{\eta}=46.79\deg$ and $\tilde{\lambda}=13.21\deg$. The ECA is the visualized size of the coverage on the Earth’s surface; the smaller the ECA, the more degraded the coverage performance becomes. Therefore, a simulation with only $\overline{\lambda}$, $\eta_{Max}$, or $\overline{\varepsilon}$ is enough to analyze if the constellation satisfies the coverage requirement. Since the coverage analyses with $\overline{\lambda}$, $\eta_{Max}$, or $\overline{\varepsilon}$ show the same results but are conducted from different perspectives, the simulation with only one of the three constraints reduces the computational cost.
@@ 3.2 Voronoi Tessellation
The problem of a continuous coverage constellation can be reduced to obtaining the circumradius of three adjacent points. For a set of discretized points, the circumradius of three adjacent points can be defined. When the circumradius is smaller than a specified value, the distance from any point within the region is shorter than the specified value. References 2 and 3 introduced the satellite triad method as a coverage analysis technique for Walker constellations. The research subject of the references was a constellation of fewer than 20 satellites. This study utilizes the Delaunay triangulation method, which was first suggested in  4, to generalize the number of satellites in the constellation.
Delaunay triangulation is a computational geometry method that subdivides discretized points into triangles[9]. This algorithm defines the Delaunay criterion for constructing Delaunay conformant triangles that do not contain other points inside the circumcircles. The Voronoi diagram (VD) is a dual graph of the Delaunay triangle (DT) and is drawn by connecting the circumcenters of the Delaunay triangles. Voronoi tessellation refers to the tiling of a plane or sphere using Voronoi diagrams. If the tiled region is a restricted closed area on the sphere, the Voronoi diagrams have boundaries cut by the region and are called bounded Voronoi diagrams (BVD) [10, 4]. The constellation coverage problem can be formulated as a Voronoi tessellation problem. The subsatellite points on the Earth’s surface are discretized points on a three-dimensional spherical surface. The area of interest is not the entire globe, and the Voronoi diagram is bounded within the target region. The spherical Delaunay triangles and spherical bounded Voronoi diagrams derive the solution; however, the word ‘spherical’ is omitted for brevity from here on.
@@
Figure 4 depicts examples of the Delaunay triangle, Voronoi diagram, and bounded Voronoi diagram. In Figure 4a, the six DTs contain $p_{k}$ as their vertices for an arbitrary subsatellite point $p_{k}$, where $k=1,2,...,T$. ${DT}_{k,l}$ are depicted as red triangles, and the circumcenters $C_{k,l}$ are blue dots, where $l=1,2,...,N_{k}$. $N_{k}$ is the number of triangles that have vertices $p_{k}$ and can be different for each $p_{k}$. Any subsatellite point possesses only one VD, and $VD_{k}$ is drawn by connecting the blue dots to the circumcenters $\textbf{C}_{k}=\left\{C_{k,l}\mid l=1,2,...,N_{k}\right\}$, as shown in Figure 4b.
The Voronoi diagram in Figure 4b can be used to solve the global coverage problem. However, for the regional coverage problem, the Voronoi diagram must be bounded within the area of interest. In particular, for the regional continuous coverage problem, the Voronoi diagram bounded within the latitude range of the AoI as a circular band can efficiently analyzes the continuity of the constellation[4]. Considering the northern area of interest, the bounded Voronoi diagram appears in Figure 4c. Then, $p_{k}$ has ${BVD}_{k}$ with different vertices $\overline{\textbf{C}}_{k}=\left\{\overline{C}_{k,\overline{l}}\mid\overline{l}=1,2,...,\overline{N}_{k}\right\}$.
The angular distance $\theta_{k,\overline{l}}$ is defined as the angular distance between the subsatellite point $p_{k}$ and the vertices $\overline{C}_{k,\overline{l}}$. Then, the maximum distance ${BVD}_{k}$ of the $k$th satellite is expressed as:
@@
Consequently, the maximum angular distance of the entire constellation is obtained as.
@@
Assuming a homogeneous constellation, the coverage performances of all satellites are the same and can be expressed as $\lambda^{*}$. Then, the continuous coverage problem statement is defined as
@@
For the global coverage problem, the maximum angular distance should be defined from $\theta_{k,l}$ and $C_{k,l}$, and the remaining procedures are the same.
@@ 3.3 APC Decomposition
The reference 7 developed the APC decomposition based on the circular convolution phenomenon between a seed satellite’s access profile, a constellation pattern vector, and a coverage timeline. The access profile between the $k$th satellite and the $j$th target point ($\textbf{{v}}_{k,j}$) defines its elements as follows:
@@
where $\tau$ is the discretized time variable and $\mathcal{J}$ is the set of target points.
The coverage timeline of constellation $\textbf{{b}}_{j}$ is derived as the summation of all access profiles as follows:
@@
The circular convolution phenomenon expresses the coverage timeline $\textbf{{b}}_{j}$ with respect to the seed satellite access profile $\textbf{{v}}_{0,j}$ and the coverage pattern vector x as
@@
This circular convolution operation can be described in a linearized form as follows:
@@
In summary, the coverage timeline of the entire constellation is obtained from the circular convolution of the seed satellite access profile and the constellation pattern vector. This concept of APC decomposition reduces the optimal CGT constellation design problem to a constellation pattern vector optimization problem.
@@ 4 Relative Motion in Adjacent Orbital Planes
Since the Walker-Delta constellation satellites are designed to have the same altitude and inclination, the relative motion between ${SAT}_{m+1,n}$ and ${SAT}_{m,n}$ can be described analytically [11]. The minimum and maximum relative angular distances ($\theta_{min}$ and $\theta_{max}$) are formulated as follows:
@@
The spherical triangle in Figure 5 derives the geometric relationship between the relative phase $\phi_{R}$ and the angular distances $\phi_{m}$ and $\phi_{m+1}$ as follows:
@@
where the differential angular distance $\phi_{m+1}-\phi_{m}$ is the relative argument of the latitude $\Delta u$ in the Walker-Delta constellation. The formula to calculate $\phi_{m+1}$ is
@@
As a result, Eq. (51) provides the argument of $\phi_{R}$ as follows:
@@
The inter-satellite link (ISL) constrains the range of relative motion so that the signal is not interfered within the link margin. Therefore, the minimum and maximum relative distances in Eqs. (46) and (47) within the specified range guarantee a smooth ISL communication.
The relative motion of the adjacent orbital plane is described in Eqs. (48), (50), and (51) and explains the relative motion between ${SAT}_{m,n}$ and ${SAT}_{m+1,n}$ where $m=1,...,P-1$ and $n=1,...,S$. When $m=P$, the relative motion between ${SAT}_{P,n}$ and ${SAT}_{1,n+F}$ is proven to be the same as the one between ${SAT}_{m,n}$ and ${SAT}_{m+1,n}$ by Eqs. (1), (2), (48), (50) and (51). As a result, if Eqs. (46) and (47) satisfy the constraint on the ISL link, then all ISL links are connected without any isolated links.
@@ 5.1 Continuous Coverage Analysis
The constellation is designed using the three constellation design methods introduced in the previous sections: Walker-Delta constellation, quasi-symmetric CGT, and BILP CGT constellations. The seed satellites for the three constellations are designed to have the repetition period of $\nu=14/1$. The inclination is 42 $\deg$ which is 3 $\deg$ –5 $\deg$ higher than the area of interest[12, 13]. The minimum elevation angle $\varepsilon_{min}$ is 15 $\deg$ and the target point is located in Seoul. The Walker-Delta constellation is analyzed assuming twobody motion, and the target area is a circle with a radius of 100 km around Seoul. The CGT constellations assume $J_{2}$ perturbation and a single target point. The sampling time or time step $t_{step}$ is 1 and 300 s for the Walker-Delta and CGT constellations, respectively, and the simulation time horizon is 1 day for both.
@@ 5.1.1 Walker-Delta Constellation
Figure 6 depicts the bounded Voronoi diagram simulation result of the Walker-Delta constellation. The global search of a smaller number of satellites obtains that 40 satellites at an inclination of 42 $\deg$ is the minimum number of satellites required to achieve the continuous coverage using the Walker-Delta constellation. The empty circles represent the $T/P/F$ parameters with duplicate positions in Eq. (14) and are precluded from the coverage analysis. The empty diamond markers indicate that the parameters did not achieve the continuous coverage. The continuous coverage solution is $i$: $T/P(S)/F$ = $42$: $40/40(1)/30$ and denoted as the black diamond. The ECA of this solution ($\lambda^{*}$) is $15.86\deg$.
5.1.2 Common Ground-track Constellation
@@
The CGT constellation’s pattern reveals its characteristics in the $\left(\Omega,M\right)$ space, such as period ratio and symmetricity (Figure 7). The gradient of the admissible set is $-N_{P}/N_{D}$ and is equal to $-\nu$ by Eqs. (25) and (15). The constellation pattern vectors are laid on the points along the admissible set. The quasi-symmetric set configures symmetrically in Figure 7 as $\textbf{{x}}_{qs}$ in the second panel of Figure 8a is equally spaced. Because $N_{qs}$ is 32 and $L$ is 288, $L/N_{qs}$ is divided into an integer $9$, and the spacing is perfectly symmetric. On the other hand, the BILP constellation pattern vector is irregularly spaced in Figures 7 and 8b. However, the BILP constellation achieves the smaller number of satellites $N_{bilp}$ as 31, while both constellations exhibit the single-fold coverage.
@@ 5.2.1 Walker-Delta Constellation
The minimum and maximum relative distances are the line distances calculated from the angular distances in Eqs. (46) and (47) (Figure 9). The upper and lower bounds of the bars indicate the relative distance ranges in the adjacent orbital planes. The colors of the bars distinguish the number of planes, and the $x$ axis represents the $F$ numbers. As this graph shows the relative motion at a glance, it is a useful tool for ISL connectivity analysis. The continuous coverage solution $i$: $T/P(S)/F$ = $42$: $40/40(1)/30$ has a relative motion range from $9559.77$ km to $9589.64$ km.